OpenStudy (anonymous):

Verify that f and g are inverse functions. f(x)=5x^2-2, x>=0, g(x)=(x+2/5)^1/2

OpenStudy (gorv):

let f(x)=y=5x^2-2 now for finding inverse first find value of x in term of y and after that switch x and y

OpenStudy (gorv):

y=5x^2-2 y+2=5x^2 (y+2)/5=x^1/2 \[x=\sqrt{\frac{ y+2 }{ 5 }}\]

OpenStudy (gorv):

now switch x and y \[y=\sqrt{\frac{ x+2 }{ 5 }}\] this is inverse of f(x) or it is g(x) hence g(x) is inverse of f(x)

OpenStudy (anonymous):

okay got it

OpenStudy (anonymous):

then for the g(x)?

OpenStudy (anonymous):

what would that be = to?

OpenStudy (anonymous):

what is the x>=0 all about..

OpenStudy (gorv):

g(x)=y lol

OpenStudy (gorv):

if we not consider x>=0 then before sqrt + and - sign will be there but x>=0 so our ans will be positive so remove negative sign

OpenStudy (gorv):

similarly find for g(x)